OpenStudy (anonymous):

For 2x – 3y = –7, when x = 4, What is the value of y?

OpenStudy (apoorvk):

Plug in x=4 in the equation, and calculate the value of y from the resulting equation.

Parth (parthkohli):

\(\Large \color{MidnightBlue}{\Rightarrow 2(4) - 3y = -7 }\) What do you think?

OpenStudy (redprince):

5 is the answers!

OpenStudy (anonymous):

5 or -5

OpenStudy (redprince):

5 as: \[2x-3y=-7\] for x=4 \[2(4)-3y=-7\] \[8-3y=-7\] \[-3y=-8-7\] \[-3y=15\] \[-3y/-3=-15/-3\] \[y=5\] get it

OpenStudy (anonymous):

thank u